Assuming that a different air mass has a relative humidity of 40% at a temperature of 25 degrees Celsius, what is the actual humidity content, expressed as g/kg?
1
Expert's answer
2020-10-05T06:43:27-0400

The relative humidity is defined by the following expression (http://www.classzone.com/vpg_ebooks/ml_earthscience_na/accessibility/ml_earthscience_na/page_392.pdf):


"RH = \dfrac{Actual\space humidity}{Maximum \space humidity}\times 100\%"

where "Maximum \space humidity" is the maximum gramms of water that one kilogram of air can hold at given temperature (units "g\/kg" ). For temperature "25\degree C", have (see https://www.engineeringtoolbox.com/humidity-ratio-air-d_686.html):


"Maximum \space humidity = 19.8\space g/kg"

Thus, the "Actual\space humidity" is:

"Actual\space humidity = \dfrac{RH\times Maximum \space humidity}{100\%} = \dfrac{40\%\times 19.8}{100\%}=\\
=7.92\space g/kg"

Answer. "7.92\space g/kg".